# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ 4. Splitting of time-reversal symmetric projections
Consider two time-reversal symmetric projection-valued maps $P,Q:\mathbb{T}^{2}\mapsto\operatorname{Proj}_{2n}(\mathcal{H})$ such that $P(k)Q(k)\equiv 0$: then $\delta(P)\delta(Q)=\delta(P+Q)$.
Proof.
We can use the previous theorem to find two decompositions $P(k)=P^{+}(k)\oplus P^{-}(k)$, $Q(k)=Q^{+}(k)\oplus Q^{-}(k)$. Then it is obvious that we can find a decomposition of $P(k)\oplus Q(k)$ as:
$P(k)\oplus Q(k)=[P^{-}(k)\oplus Q^{-}(k)]\oplus[P^{+}(k)\oplus Q^{+}(k)]$
Now we can use Lemma 2.6, namely $\operatorname{Ch}(P^{-}\oplus Q^{-})=\operatorname{Ch}(P^{-})+\operatorname{Ch}(Q^{-})$. So we can conclude using the previous theorem:
$\delta(P\oplus Q)=e^{i\pi\operatorname{Ch}(P^{-}\oplus Q^{-})}=e^{i\pi\operatorname{Ch}(P^{-})}\cdot e^{i\pi\operatorname{Ch}(Q^{-})}=\delta(P)\cdot\delta(Q).\qed$
5. Pseudo-periodic frames
One approach to the topological study of projection-valued maps is to investigate under which assumption is it possible to build a periodic and symmetric frame for a given projection-valued map. To be precise, given $P:\mathbb{T}^{d}\to\operatorname{Proj}_{n}(\mathcal{H})$, the goal is to find a periodic and continuous (or analytic) collection of orthonormal vectors $\{v_{j}(k)\}_{j\in 1,\cdots,n}$ that span $\operatorname{Ran}(P(k))$ for all $k\in\mathbb{T}^{d}$. In some circumstances, this operation turns out to be impossible, and non trivial topological invariants like $\operatorname{Ch}$ and $\delta$ are seen as topological obstructions in the construction of a frame [panati2007triviality, monaco2015symmetry, fiorenza2016z].
More recently, a more refined goal has been researched, namely how close one can get to a global, symmetric, continuous, and periodic frame when its construction is topologically obstructed: see e.g. [cornean2019parseval, auckley2018parseval]. One relaxes the notion of periodicity to one of pseudo-periodicity (or periodicity up to a phase) for the minimal number possible of vectors in the basis. For example, for a non-symmetric projection-valued map the following result from [monaco2023topology] holds.
Theorem 5.1 (non-symmetric pseudo-periodic frame).
Consider a projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$. Then it is always possible to construct a pseudo-periodic frame $\{v_{j}(k)\}_{j\in\{1,\cdots,n\}}$ such that all vectors are periodic in $k$ except the first that satisfies
$v_{1}(\pi,k_{2})=e^{i\operatorname{Ch}(P)k_{2}}v_{1}(-\pi,k_{2}),\quad v_{1}(k_{1},0)=v_{1}(k_{1},2\pi)\quad\forall k_{1},k_{2}\in\mathbb{T}^{2}.$
Proof.
If $\{v_{j}\}_{j\in\{1,\cdots,n\}}$ is an orthonormal basis of $P(0,0)$, then $\{u_{j}(t,k_{2})=U(t,k_{2})v_{j}\}_{j\in\{1,\cdots,n\}}$ constitutes an orthonormal basis of $P(t,k_{2})$ that depends continuously in $t$ and $k_{2}$. This is periodic in $k_{2}$ but in general will not be periodic in $t$. The aperiodicity is once again measured by the matching matrix $\alpha(k_{2})\in U(n)$, with:
$u_{j}(\pi,k_{2})=\sum_{a=1}^{n}\alpha_{j,a}(k_{2})u_{j}(-\pi,k_{2}).$
Since $\det(\alpha(k_{2}))]=\operatorname{Ch}(P)$ the matrix:
$\beta(k_{2})=\begin{pmatrix}e^{(i\operatorname{Ch}(P)k_{2})}&0\\ 0&\mathds{1}_{n-1}\end{pmatrix}\alpha(k_{2})^{-1}$
is such that $\det(\beta(k_{2}))]=0$. So, Using Theorem A.7, there is a homotopy $\beta_{s}(k_{2})$ with $\beta_{1}(k_{2})=\beta(k_{2})$ and $\beta_{0}(k_{2})=\mathds{1}$. Therefore, the vectors
$v_{j}(t,k_{2})=\sum_{a=1}^{n}\left[\beta_{(t+\pi)/2\pi}(k_{2})\right]_{j,a}u_{j}(t,k_{2})\quad\mbox{for}\quad j\in\{1,\cdots,n\}$
constitute an orthonormal basis of $\operatorname{Ran}(P(t,k_{2}))$ and are such that:
$v_{j}(\pi,k_{2})$ $=\sum_{a=1}^{n}[\beta(k_{2})]_{j,a}u_{a}(\pi,k_{2})=\sum_{a,b=1}^{n}[\beta(k_{2})]_{j,a}\alpha_{a,b}(k_{2})u_{b}(-\pi,k_{2})=$
$=\sum_{a=1}^{n}[\beta(k_{2})\alpha(k_{2})]_{j,a}u_{a}(-\pi,k_{2})=\begin{cases}e^{i\operatorname{Ch}(P)k_{2}}v_{1}(-\pi,k_{2})&\mbox{for }j=1,\\ v_{j}(-\pi,k_{2})&\mbox{otherwise}.\end{cases}\qed$
Instead, in presence of a time-reversal symmetry $T$ such that $T^{2}=-\mathds{1}$, the most natural symmetric constraint we can impose is that the frame must be composed by pairwise symmetric vectors $v(k),u(k)$ such that Tv(k)=u(k)Tv(k)=u(-k). This is the optimal condition because it is not possible to relate a vector to itself by imposing Tv(k)=v(k)Tv(k)=v(-k). In fact, it would lead to $Tv(0)=v(0)$ and consequently to $v(0)=T^{2}v(0)=v(0)$ which is absurd. This relation can instead be imposed when $T^{2}=\mathds{1}$, as discussed in [cornean2016construction] and [fiorenza2016z].
Theorem 5.2 (Pseudo-periodic symmetric frame).
Consider a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$. Then, it is always possible to construct a symmetric and pseudo-periodic frame
$\{v_{1},\cdots,v_{n}(k),-Tv_{1}(-k),\cdots,-Tv_{n}(-k)\}$
of $P(k)$ such that, if $\delta(P)=1$, the frame has full periodicity, while if $\delta(P)=-1$, the frame is periodic with the only exceptions being:
$v_{1}(\pi,k_{2})=e^{ik_{2}}v_{1}(-\pi,k_{2})\quad\text{and}\quad v_{n+1}(\pi,k_{2})=e^{-ik_{2}}v_{n+1}(-\pi,k_{2})\quad\forall k_{2}\in\mathbb{T}^{1}.$
Proof.
The construction of the symmetric frame can be done by applying Theorem 4.2: if $\delta(P)=1$ we can build a decomposition with any even Chern number, in particular with a vanishing one. Instead, if $\delta(P)=-1$ we can build a decomposition $P=P^{-}\oplus P^{+}$ with $P^{-}$ having any odd Chern number, in particular with $\operatorname{Ch}(P^{-})=1$. To conclude, we need to apply Theorem 5.1 to $P^{-}$ and then symmetrize the frame obtained. ∎
Remark 5.3 (Symmetry vs periodicity).
Theorems 5.1 and 5.2 together with Remark 4.1 yield an interesting fact. If we have a symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ we have that $\operatorname{Ch}(P)=0$. So, it is possible to construct a periodic and continuous frame. However, if we are looking for a frame that respects the symmetry $T$, it is not possible to construct a continuous, symmetric, and periodic frame when $\delta(P)=-1$. This means that there is a natural question that arises: what is the natural conflict between symmetry and periodicity? The answer follows naturally from Theorem 5.2. In fact, if we look at the non-periodic vectors $v_{1},v_{n+1}$ we have that they span a time-reversal symmetric projection-valued map $\tilde{P}:\mathbb{T}^{2}\to\operatorname{Proj}_{2}(\mathcal{H})$ such that $\delta(\tilde{(}P))=-1$ and that contains the non-trivial topological information of $P$. Instead, the other periodic vectors $v_{2},\cdots,v_{n-1},v_{n+2},\cdots,v_{2n}$ span a time-reversal symmetric projection-valued map $\hat{P}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n-2}(\mathcal{H})$ such that $\delta(\hat{P})=1$. So if one looks for periodicity over symmetry it is possible to construct a frame in which only two vectors are not pairwise symmetric while the others are perfectly continuous, symmetric and periodic.
Following this lines, by repeating the construction of [cornean2019parseval], one can use the same “Hilbert space doubling” trick employed there to the map $\tilde{P}$ in order to span it with four periodic, non-orthogonal vectors, which however come in time-reversal-symmetry-related Kramers pair. Together with the symmetric and periodic frame for $\hat{P}$, these $m+2$ vector constitute a continuous, symmetric and periodic Parseval frame spanning the projection $P$.
6. Homotopy problem
In Definition 1.4 we gave a notion of symmetric unitary equivalence and a notion of symmetric homotopy. The goal of this section is to prove that $\delta$ is the only topological information needed in order to connect two different projection-valued maps via symmetric homotopies or symmetric unitary equivalences. Before proving this result, we need a technical lemma that essentially allows us to construct symmetric, continuous, and periodic frames for infinite-dimensional time-reversal symmetric projection-valued maps.
Lemma 6.1.
@@ 6. Homotopy problem
If $\mathcal{H}$ is a separable Hilbert space with $\dim(\mathcal{H})=\infty$ and $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ is a continuous time-reversal symmetric projection-valued map with TP(k)=P(k)TTP(k)=P(-k)T, then there is always a continuous and periodic collection of orthonormal vectors $\{v_{j}(k)\}_{j\in\mathbb{N}}$ such that
@@
We can start by using Lemma A.1 to select a discrete quaternionic basis $\{v_{j}\}_{j\in\mathbb{N}}$ of $\ker(P(0,0))$. This basis, together with $\{Tv_{j}\}_{j\in\mathbb{N}}$, is a complex basis of $\ker(P(0,0))$. Then we can use Proposition 2.3 to construct a unitary-valued map $U(t,k_{2})$ such that
@@
We can construct an orthonormal basis of $\ker(P(t,k_{2}))$ using the vectors
@@
and $\mathcal{K}$ is the standard complex conjugation. The aperiodicity is expressed by the fact that
@@
$\begin{split}\mathcal{K}J^{-1}I(\pi,0)I(-\pi,0)^{-1}&=I(-\pi,0)T^{-1}I(-\pi,0)^{1}=I(-\pi,0)I(\pi,0)^{-1}\mathcal{K}J^{-1}=\\ &=(I(\pi,0)I(-\pi,0)^{-1})^{-1}\mathcal{K}J^{-1}.\end{split}$
@@
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J=\\ &=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\mathcal{K}J^{-1}I(-\pi,k_{2})I(\pi,k_{2})^{-1}\beta(k_{2})\mathcal{K}J\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
@@
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}I(\pi,k_{2})I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\mathcal{K}J^{-1}I(-\pi,-k_{2})I(\pi,-k_{2})^{-1}\beta(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
@@
Then $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ is the collection of vectors we were looking for, as it is periodic and continuous because so is $\tilde{I}$. Moreover, elements such as $(0,\delta_{j})$ together with $(\delta_{j},0)$ form an orthonormal basis of $l^{2}(\mathbb{N})^{2}$, so their images $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ together with $\{\tilde{u}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(0,\delta_{j})\}_{j\in\mathbb{N}}$ form a discrete orthonormal basis of $\ker(P(t,k_{2}))$. Most importantly, the relation $(\delta_{j},0)=\mathcal{K}J(0,\delta_{j})$ implies that $\tilde{v}_{j}(t,k_{2})=T\tilde{u}_{j}(-t,-k_{2})$ for all $j\in\mathbb{N},t\in[-\pi,\pi],k_{2}\in\mathbb{T}^{1}$. So, the vectors
@@
Two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are symmetrically unitarily equivalent and homotopic if and only if $\delta(P_{0})=\delta(P_{1})$.
@@
We already know one part of the statement: in fact, Proposition 3.3 states that if two maps are unitarily equivalent or homotopic, then their $\delta$ invariants are equal. So, we want to construct a symmetric unitary equivalence between $P_{0}$ and $P_{1}$ under the assumption that $\delta(P_{0})=\delta(P_{1})$. We can start by applying Theorem 5.2 to construct two (pseudo-periodic) frames
@@
where the added vectors span respectively $\ker(P_{0}(k))$ and $\ker(P_{1}(k))$ and have full periodicity with exceptions
@@
where the added vectors have full periodicity and are two orthonormal bases of $\ker(P_{0}(k))$ and $\ker(P_{1}(k))$, respectively.
@@
Now we want to solve the homotopy problem, so if $\delta(P_{0})=\delta(P_{1})$, we want to exhibit a symmetric homotopy between the two projection-valued maps. From the previous point we can create a symmetric unitary equivalence $P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}$ with TV(k)=V(k)TTV(k)=V(-k)T. The first step is to define a unitary-valued map $V_{0}:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that $TV_{0}(k)=V_{0}(-k)T$ and $P_{0}(k)=V_{0}(k)P_{0}(k)V_{0}(k)^{-1}$. If $\dim(\mathcal{H})=\infty$ we can define $V_{0}(k)\equiv\mathds{1}$. Instead, if $\dim(\mathcal{H})=2N<\infty$, we need to do something more. First, we need to use lemma A.1 to select a basis on which $T$ acts as $J\mathcal{K}$. Then we need to observe that the condition TV(k)=V(k)TTV(k)=V(-k)T means that $J\overline{V(k)}=V(-k)J$. Set $\lambda(k):=\det(V(k))$, and consider the triangle $\triangle$ inside $\pi,\pi]^{2}$ with vertices
@@
Then on the points $k_{\star}$ fixed by the involution $k\mapsto-k$ we have that $J=V(k_{\star})^{t}JV(k_{\star})$, so $V(k_{\star})\in\mathrm{Sp}(2n)$. Moreover, on the segments $\overline{Q_{1}Q_{2}},\overline{Q_{1},Q_{3}},\overline{Q_{2}Q_{3}}$, we have that $\lambda(k)=\overline{\lambda(-k)}$. Therefore, in those segments, we can apply Lemma A.3 to have that
@@
Since the map is defined also inside the triangle, the winding number of $\lambda$ in $\triangle$ must be zero, so $2m=2l+2h$. Now we can use the quasi-periodic frame $\{v_{1}^{0}(k),\cdots,-Tv_{N}^{0}(-k)\}$ to define the unitary operator $V_{0}(k)$ with:
@@
$\begin{split}V_{0}(-\pi,k_{2})v_{1}^{0}(\pi,k_{2})&=V_{0}(-\pi,k_{2})e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i\pi h(\delta)}e^{i(-m\pi+lk_{2})}v_{1}^{0}(-\pi,k_{2})=\\ &=e^{-2im\pi}e^{i(m\pi+lk_{2})}e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i(m\pi+lk_{2})}v_{1}^{0}(\pi,k_{2})=\\ &=V_{0}(\pi,k_{2})v_{1}^{0}(\pi,k_{2})\end{split}$
In both cases $\{\dim(\mathcal{H})=\infty,\dim(\mathcal{H})=2N\}$ we want to define a homotopy $V_{t}$ between $V_{0}$ and $V$. To do so, we need to work on $\triangle$. Notice that on the sides of the triangle $\overline{Q_{1}Q_{2}}$, $\overline{Q_{1}Q_{3}}$ and $\overline{Q_{2}Q_{3}}$ the projection-valued map behaves as a one-dimensional time-reversal symmetric projection-valued map:
@@
So we can use Theorem 2.2 to build $V_{t}(k)$ for $k\in\partial\triangle$ with $TV_{t}(k)=V_{t}(-k)T$ for all $t\in[0,1],k\in\partial\triangle$. This means that $V_{t}(k)$ is defined for $(t,k)\in\partial([0,1]\times\triangle)$. This region however, is homeomorphic to $S^{2}$ and since $\pi_{2}(U(2N))=\{0\}=\pi_{2}(\mathcal{U}(\mathcal{H}))$, thanks to Theorems A.7 and A.8, it is always possible to extend the definition inside for $(t,k)\in[0,1]\times\triangle$. As a final adjustment, we just need to define the map in the complementary triangle by imposing $V_{t}(k)=T^{-1}V_{t}(-k)T$ for $k\in[-\pi,\pi]^{2}\setminus\triangle$. Then $V_{t}(k)$ is continuous, periodic and symmetric and therefore:
@@
Remark 6.3.
We want to emphasize the role that symmetry plays in this topological study. It is known that projection-valued maps with equal Chern number are homotopic and unitarily equivalent. However, Remark 4.1 states that time-reversal symmetric projection-valued maps always have a trivial Chern number and should always be unitarily equivalent and homotopic. How can this statement agree with the previous theorem? The answer is simple: If two time-reversal symmetric projection-valued maps have different $\delta$ invariants than any homotopy or unitary equivalence connecting the two must break the symmetry constraints present in Definition 1.4. In this way our approach is consistent with the existing theory about vector bundles.

### v2

@@ 5.4. Emerging spin–type simmetry
If $\mathcal{H}$ is a separable Hilbert space with $\dim(\mathcal{H})=\infty$ and $P:\mathbb{T}^{2}\to\operatorname{Proj}_{\infty}(\mathcal{H})$ is a continuous time-reversal symmetric projection-valued map with TP(k)=P(k)TTP(k)=P(-k)T, then there is always a continuous and periodic collection of orthonormal vectors $\{v_{j}(k)\}_{j\in\mathbb{N}}$ such that
@@
We can start by using Lemma A.1 to select a discrete quaternionic basis $\{v_{j}\}_{j\in\mathbb{N}}$ of $\ker P(0,0)$. This basis, together with $\{Tv_{j}\}_{j\in\mathbb{N}}$, is a complex basis of $\ker P(0,0)$. Then we can use Proposition 2.3 to construct a unitary-valued map $U(t,k_{2})$ such that
@@
We can construct an orthonormal basis of $\ker P(t,k_{2})$ using the vectors
@@
and $\mathcal{K}$ is the standard complex conjugation. The lack of periodicity is expressed by the fact that
@@
$\begin{split}\mathcal{K}J^{-1}I(\pi,0)I(-\pi,0)^{-1}&=I(-\pi,0)T^{-1}I(-\pi,0)^{1}=I(-\pi,0)I(\pi,0)^{-1}\mathcal{K}J^{-1}\\ &=(I(\pi,0)I(-\pi,0)^{-1})^{-1}\mathcal{K}J^{-1}.\end{split}$
@@
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J\\ &=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\mathcal{K}J^{-1}I(-\pi,k_{2})I(\pi,k_{2})^{-1}\beta(k_{2})\mathcal{K}J\mathcal{K}J\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
@@
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J\\ &=I(\pi,k_{2})^{-1}I(\pi,k_{2})I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta(-k_{2})\mathcal{K}J\\ &=I(\pi,k_{2})^{-1}\mathcal{K}J^{-1}I(-\pi,-k_{2})I(\pi,-k_{2})^{-1}\beta(-k_{2})\mathcal{K}J\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
@@
Then $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ is the collection of vectors we were looking for, as it is periodic and continuous because so is $\tilde{I}$. Moreover, elements such as $(0,\delta_{j})$ together with $(\delta_{j},0)$ form an orthonormal basis of $l^{2}(\mathbb{N})^{2}$, so their images $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ together with $\{\tilde{u}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(0,\delta_{j})\}_{j\in\mathbb{N}}$ form a discrete orthonormal basis of $\ker P(t,k_{2})$. Most importantly, the relation $(\delta_{j},0)=\mathcal{K}J(0,\delta_{j})$ implies that $\tilde{v}_{j}(t,k_{2})=T\tilde{u}_{j}(-t,-k_{2})$ for all $j\in\mathbb{N},t\in[-\pi,\pi],k_{2}\in\mathbb{T}^{1}$. So, the vectors
@@
Theorem 5.6 (Emerging spin-type symmetry).
Let $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ be a time-reversal symmetric projection-valued map. Then:
  1. (1)
    there exists a spin-type symmetry, namely a continuous unitary-valued map $S:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that
    $S(k)^{2}=\mathds{1},\qquad[S(k),P(k)]=0,\qquad TS(k)=-S(-k)T,\qquad\forall k\in\mathbb{T}^{2};$
  2. (2)
    the $\mathbb{Z}_{2}$ invariant of $P$ can be computed as a Spin-Chern number with respect to any spin-type symmetry $S$:
    $\delta(P)=(-1)^{\operatorname{Ch}_{S}(P)}.$
Proof.
The existence of a map $S$ as in (1) is equivalent to the existence of symmetric splittings of both $P$ and its orthogonal complement $P^{\perp}:=\mathds{1}-P$
$P(k)=P^{+}(k)+P^{-}(k),\qquad P^{\perp}(k)=P^{\perp,+}(k)+P^{\perp,-}(k).$
Indeed, given such splittings, the operator
$S(k):=\mathds{1}-2\big(P^{+}(k)+P^{\perp,+}(k)\big)$
satisfies $S(k)^{2}=\mathds{1}$, since $P^{+}(k)$ and $P^{\perp,+}(k)$ are orthogonal, hence $P^{+}(k)+P^{\perp,+}(k)$ is a projection. Moreover, $[S(k),P(k)]=0$, since both $P^{+}(k)$ and $P^{\perp,+}(k)$ commute with $P(k)$. Finally,
$TS(k)$ $=T\big(\mathds{1}-2P^{+}(k)-2P^{\perp,+}(k)\big)$
$=\big(\mathds{1}-2P^{-}(-k)-2P^{\perp,-}(-k)\big)T=-S(-k)T,$
where the last equality follows from the identity
$P^{+}(k)+P^{\perp,+}(k)+P^{-}(k)+P^{\perp,-}(k)=\mathds{1}.$
A symmetric splitting of $P$ can always be obtained by means of Theorem 5.1. The existence of such a splitting for $P^{\perp}$ follows from the same result when $\mathcal{H}$ is finite-dimensional, since in this case $P^{\perp}$ is a finite-rank time-reversal symmetric projection-valued map.
In the infinite-dimensional case, Lemma 5.5 guarantees the existence of a family $\{v_{i}\}_{i\in\mathbb{N}}$ of continuous vector-valued functions such that $\{v_{i}\}_{i\in\mathbb{N}}\sqcup\{Tv_{i}\}_{i\in\mathbb{N}}$ defines a frame for $P^{\perp}$. One then obtains the desired symmetric splitting of $P^{\perp}$ by defining
$P^{\perp,+}(k):=\sum_{i=0}^{\infty}|v_{i}(k)\rangle\langle v_{i}(k)|,\qquad P^{\perp,-}(k):=\sum_{i=0}^{\infty}|Tv_{i}(k)\rangle\langle Tv_{i}(k)|.$
Statement (2) follows directly from formula (11) and from Theorem 5.1. Indeed, any spin-type symmetry $S$ induces a symmetric splitting $P=P^{+}+P^{-}$ by setting $P^{\pm}:=P(S\pm\mathds{1})/2$, and therefore
$\delta(P)=(-1)^{\operatorname{Ch}(P^{\pm})}=(-1)^{\operatorname{Ch}_{S}(P)}.\qed$
5.5. Completeness of the invariant
In Definition 1.8 we introduced a notion of Murray–von Neumann equivalence, a notion of symmetric unitary equivalence, and a notion of symmetric homotopy for time-reversal symmetric projectoion-valued maps. The goal of this section is to prove that $\delta$ is the only topological information needed to connect two different projection-valued maps by means of a symmetric homotopy. This is sufficient to show the completeness of the invariant with respect to all three equivalence relations, since homotopy equivalence is the weakest among them (see Remark 1.9).
This result is partly known for other formulations of the invariant: for instance, completeness with respect to partial isometries (Murray–von Neumann equivalence) is discussed in the last section of [fiorenza2016z], while in [Cornean_2017] completeness with respect to homotopy is discussed for the family of matching matrices, although an explicit homotopy between the associated projection-valued maps is not constructed — precisely what we will instead carry out in the following proof.
@@ 5.5. Completeness of the invariant
Two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are symmetrically Murray–von Neumann equivalent, unitarily equivalent and homotopic if and only if $\delta(P_{0})=\delta(P_{1})$.
@@
We already know one part of the statement: in fact, Proposition 5.2 states that if two maps are Murray–von Neumann equivalent, unitarily equivalent or homotopic, then their $\delta$ invariants are equal. So, we want to construct a symmetric unitary equivalence between $P_{0}$ and $P_{1}$ under the assumption that $\delta(P_{0})=\delta(P_{1})$; the restriction of this unitary to the range of $P_{0}$ would then yield a Murray–von Neumann equivalence. We can start by applying Theorem 5.1 to construct two pseudo-periodic symmetric frames
@@
where the added vectors span respectively $\ker P_{0}(k)$ and $\ker P_{1}(k)$ and have full periodicity with exceptions
@@
where the added vectors have full periodicity and are two orthonormal bases of $\ker P_{0}(k)$ and $\ker P_{1}(k)$, respectively.
@@
Now we pass to consider homotopy equivalence, so if $\delta(P_{0})=\delta(P_{1})$, we want to exhibit a symmetric homotopy between the two projection-valued maps. By the previous discussion we can create a symmetric unitary equivalence $P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}$ with TV(k)=V(k)TTV(k)=V(-k)T. The first step is to define a unitary-valued map $V_{0}:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that $TV_{0}(k)=V_{0}(-k)T$ and $P_{0}(k)=V_{0}(k)P_{0}(k)V_{0}(k)^{-1}$. If $\dim(\mathcal{H})=\infty$ we can define $V_{0}(k)\equiv\mathds{1}$. Instead, if $\dim(\mathcal{H})=2N<\infty$, we need to do something more. First, we need to use lemma A.1 to select a basis on which $T$ acts as $J\mathcal{K}$. Then we need to observe that the condition TV(k)=V(k)TTV(k)=V(-k)T means that $J\overline{V(k)}=V(-k)J$. Set $\lambda(k):=\det V(k)$, and consider the triangle $\triangle$ inside $\pi,\pi]^{2}$ with vertices
@@
Then on the points $k_{\star}$ fixed by the involution $k\mapsto-k$ we have that $J=V(k_{\star})^{t}JV(k_{\star})$, so $V(k_{\star})\in\mathrm{Sp}(2n)$. Moreover, on the segments $\overline{Q_{1}Q_{2}},\overline{Q_{1},Q_{3}},\overline{Q_{2}Q_{3}}$, we have that $\lambda(k)=\overline{\lambda(-k)}$. Therefore, on those segments, we can apply Lemma A.3 to have that
@@
Since the map is defined also inside the triangle, the winding number of $\lambda$ in $\triangle$ must be zero, so $2m=2l+2h$. Now we can use the pseudo-periodic frame $\{v_{1}^{0}(k),\cdots,-Tv_{N}^{0}(-k)\}$ to define the unitary operator $V_{0}(k)$ with:
@@
$\begin{split}V_{0}(-\pi,k_{2})v_{1}^{0}(\pi,k_{2})&=V_{0}(-\pi,k_{2})e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i\pi h(\delta)}e^{i(-m\pi+lk_{2})}v_{1}^{0}(-\pi,k_{2})\\ &=e^{-2im\pi}e^{i(m\pi+lk_{2})}e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i(m\pi+lk_{2})}v_{1}^{0}(\pi,k_{2})\\ &=V_{0}(\pi,k_{2})v_{1}^{0}(\pi,k_{2})\end{split}$
In both cases $\{\dim(\mathcal{H})=\infty,\dim(\mathcal{H})=2N\}$ we want to define a homotopy $V_{t}$ between $V_{0}$ and $V$. To do so, we need to work on $\triangle$. Notice that on the sides of the triangle $\overline{Q_{1}Q_{2}}$, $\overline{Q_{1}Q_{3}}$ and $\overline{Q_{2}Q_{3}}$ the projection-valued map behaves as one-dimensional time-reversal symmetric projection-valued maps:
@@
So we can use Theorem 2.2 to build $V_{t}(k)$ for $k\in\partial\triangle$ with $TV_{t}(k)=V_{t}(-k)T$ for all $t\in[0,1],k\in\partial\triangle$. This means that $V_{t}(k)$ is defined for $(t,k)\in\partial([0,1]\times\triangle)$. This region however, is homeomorphic to $S^{2}$ and since $\pi_{2}(U(2N))=\{0\}=\pi_{2}(\mathcal{U}(\mathcal{H}))$, thanks to Theorems A.6 and A.7, it is always possible to extend the definition inside for $(t,k)\in[0,1]\times\triangle$. As a final adjustment, we just need to define the map in the complementary triangle by imposing $V_{t}(k)=T^{-1}V_{t}(-k)T$ for $k\in[-\pi,\pi]^{2}\setminus\triangle$. Then $V_{t}(k)$ is continuous, periodic and symmetric and therefore: